# Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer

### v2

@@ II Results
Figure 4: Experiments: Chemistry on large basis sets. (a) 58 qubits are used to model the N2 dissociation (cc-pVDZ basis set). (b) 45 qubits are used for the [2Fe-2S] cluster (TZP-DKH basis set) and (c) 77 qubits for the [4Fe-4S] cluster (TZP-DKH basis set). The top panels show a 3-dimensional representation of the geometry of each molecule. The middle panels show the qubits selected on a Heron quantum processor layout, following the same color convention as panel (b) in Fig. 3. The bottom panel in (a) shows the potential energy surface comparison, as well as the energy difference $\Delta E$ between the Heat-Bath Configuration Interaction (HCI) energy and the energies obtained from different methods. The brown scatter plot shows the value of $E^{(k)}$ for all batches of configurations and the connected dots show $\min_{k}(E^{(k)})$. The bottom panel in (b) shows the energy-variance analysis for three different eigenstates that both HCI and our method find upon increasing the value of $d$, as labeled by the color bar. For each approximate eigenstate $|\psi^{(k)}\rangle$, the horizontal axis $\Delta H=\langle\psi^{(k)}|\hat{H}^{2}|\psi^{(k)}\rangle-\langle\psi^{(k)}|\hat{H}|\psi^{(k)}\rangle^{2}$. The bottom panel in (c) shows a comparison of the energy-variance analysis applied to quantum measurement outcomes and bitstrings (with the correct particle number) sampled from the uniform distribution. The DMRG energy in panels (b) and (c) is from Ref. [45].
@@
Iron–sulfur (FeS) clusters are molecular ensembles of sulfide-linked 1- to 8-iron centers in variable oxidation states. They are important cofactors in biological processes ranging from nitrogen fixation to photosynthesis and respiration [46]. Their electronic structure, with multiple low-lying states of differing electronic and magnetic character, is responsible for their rich chemistry. At the same time, they pose considerable challenges for experimental studies and numerical tools. For our experiments we consider the synthetic [Fe2S2(SCH3)4]-2 cluster [4], abbreviated [2Fe-2S] and used in numerical studies to mimic the oxidized dimers prominently found in ferredoxins [47].
@@
The qubit mapping of the LUCJ circuit on the Heron processor for [2Fe-2S] is shown in the middle panel of Fig. 4 (b). We consider $K=10$ batches of configurations and $|\tilde{\mathcal{X}}|=2.4576\cdot 10^{6}$ measurement outcomes. For the [2Fe-2S] cluster, we perform an energy-variance analysis of the low-energy spectrum of the molecule. The energy-variance analysis is a tool routinely used in classical computational electronic structure to capture the convergence of the approximate eigenstate energy for different levels of accuracy of a computational method [48]. Here we use energy-variance analysis to assess the convergence as a function of $d$, which is directly related to quantum and classical accuracy, runtimes, and costs. If one can statistically sample from a good approximation of an eigenstate, points at finite number of samples will be distributed linearly in the energy-variance plane [48]. This gives us a tool to detect eigenstates for both quantum and classical methods.
The bottom panel in Fig. 4(b) shows an energy-variance comparison of HCI and our estimator. The accuracy of both estimators is quantified by the number of electronic configurations $d$ that span the subspace for diagonalization. For $d>250\cdot 10^{3}$, we warm start with the ${\bf{n}}$ obtained from the method at $d=250\cdot 10^{3}$, and only two recovery iterations are considered after. As we increase $d$, three eigenstates appear, which we label A, B, and C. At lower values of $d$, the data follows the expected linear relation for eigenstate A. Beyond a critical value of $d$, both the energy and variance change abruptly, discovering another set of points following a linear relation, which correspond to a second eigenstate B, of lower energy. The resulting mean occupations ${\bf{n}}$ of eigenstates A and B are substantially different [29], confirming that we have found two distinct eigenstates. The third cluster of points in the energy-variance plane (Eigenstate C) was found by warm starting the procedure, for all values of $d$ considered, with the ${\bf{n}}$ obtained from the second cluster (Eigenstate B) at $d=25\cdot 10^{6}$. We then extrapolate the quantum data following the linear energy-variance relations, finding agreement between the quantum extrapolated data and the best HCI calculation performed and the Density Matrix Renormalization Group (DMRG) calculaiton from Ref. [45], showcasing the ability of our methods to perform precision many-body physics.
@@
The circuits considered for the N2 experiments and [2Fe-2S] reached sizes of approximately 1-1.5k two-qubit gates. We now test the quality of the signal in noisy circuits that test the limits of Heron devices, using up to 6400 nodes of Fugaku for the classical processing. We consider the synthetic [Fe4S4(SCH3)4]-2 cluster [4], abbreviated [4Fe-4S], a representative of nature’s cubanes, whose ground state deduction from experimental measurements was an early success of inorganic spectroscopy [49]. The LUCJ circuit employed for this molecular species contains approximately 3.5k two-qubit gates. The qubit mapping on the Heron processor for [4Fe-4S] is shown in the middle panel of Fig. 4 (c). As in the previous experiment, we consider $K=10$ batches of configurations and $|\tilde{\mathcal{X}}|=2.4576\cdot 10^{6}$ measurement outcomes. For $d>250\cdot 10^{3}$, configuration recovery is warm started with the ${\bf{n}}$ obtained from the method at $d=250\cdot 10^{3}$, and only two iterations are then performed.
@@
This last set of experiments sheds light on the quality of the quantum signal that is passed to the configuration recovery at these large circuit sizes. The bottom right panel in Fig. 4 (c) shows a comparison of the energy-variance analysis from measurement outcomes obtained from the Heron device and configurations sampled from the uniform distribution. We see that, even if the quantum solutions produced are worse than other classical methods, the energy and variance obtained from quantum data are significantly lower than those obtained from uniformly distributed configurations (i.e. pure noise), on subspaces of the same size. This confirms that there is a valuable signal at circuit sizes of approximately 3.5k two-qubit gates.
III Discussion
Current quantum computers in isolation can be used as tools to explore calculations beyond brute-force [50, 51] on problems that match the interactions of the quantum devices. The assistance of classical distributed computations enables addressing problems that go beyond this constraint, with workflows such as the one used here for chemistry.
We now discuss the implications of our results in the search for quantum advantage. Any computation is ranked against three parameters: runtime, energy or cost, and accuracy. While the first two are often easily measured, ranking by accuracy is not always straightforward. However, the class of computational methods that strictly produce upper bounds to the ground-state energy defines an unconditional accuracy metric, which can be used to assess quantum advantage. Within this class of methods, a lower energy is ranked as higher quality, with no extra information required. This, for example, allowed us to benchmark our results against classical selected configuration interaction and uniform configuration sampling on an experiment on 77 qubits, without access to exact solutions. Additionally, the approximate solutions found here can be stored in classical memory, which permits a classical prover to certify them, and allows their manipulation by further classical processing.
The estimator can be applied to simulation tasks other than quantum chemistry, if the target ground state energy can be captured by a sparse wavefunction. This motivates looking into useful quantum circuits with supports in the computational basis scaling polynomially with problem size. Although in principle other estimators, such as the standard ones used in variational quantum eigensolvers, have bounded variance for any wavefunction, the dire scaling of number of measurements to estimate energies makes them impractical for the molecules targeted in this work [3].
Reconstructing the exact ground state probability distribution on a quantum computer is not required to get accurate energy approximations, as long as one is sampling from the ground state support. It will be compelling to use these methods on future quantum devices with lower error rates and correspondingly reduced classical processing requirements.
@@ III Discussion
We have used a LUCJ class of quantum circuits that can reproduce a low-rank decomposition and sparsification of the quantum unitary CCSD (qUCCSD), which allowed us to keep circuit depths manageable  [40]. Better quantum noise rates would allow to access deeper quantum circuits, improving their approximation of the qUCCSD theory and eventually surpassing qUCCSD itself if enough parameters are given. The inclusion of more density-density interactions, takes LUCJ further away from non-interacting fermionic computations, increasing the classical simulation overhead [52]. We have performed optimization-free experiments, exploiting the connection between LUCJ and classical coupled cluster theory, yet closing a quantum-classical optimization could further improve the quality of the solutions.
@@
It is important to evaluate the methods presented here in the context of variational classical heuristics to highlight their different domains of applicability [53]. To improve on other classical CI methods, one would need to sample important configurations with a significant number of excitations over the HF state. Variational quantum Monte Carlo methods [54, 55], of paramount importance in many-body physics, are sensitive to the structures of the probability distributions they are modeling, not just to their supports. Methods based on tensor networks [56, 57] are successfully used to tackle strongly correlated problems in chemistry, granted the ability to converge their energies with bond dimensions. Convergence can be challenging in some cases [58, 59] because of its computational cost and its sensitivity to the nature and ordering of the basis-set orbitals. Characterizing the domain of applicability of different classical and quantum heuristics facilitates combining them in quantum-centric supercomputing environments, shifting the search for quantum advantage towards practical problems.
Author contributions
Design of the workflow and experiments: J. R.-M., M. M., W. K., K. S., M. T., A. M. Implementation of the workflow: J. R.-M., M. M., A. J.-A., S. M., S. S., T. S., I. S., R.-Y. S., K. J. S. Tuning and calibration of the Heron processor: H.H. Execution of the quantum part of the workflow on Heron: M. M., P. J., M. T. Execution of the classical part of the workflow on Fugaku: T. S., R.-Y. S., S. Y. Numerical benchmarks: J. R.-M., M.M., T. S., K. J. S. Analytical derivations: J. R.-M., M.M., W. K., K. S., M. T., A. M. All authors contributed to the manuscript writing and data analysis.
Acknowledgements.
We thank the IBM Quantum Service and Data team for help with the workflow execution. The authors acknowledge feedback and insightful conversations with Alberto Baiardi, Sergey Bravyi, Giuseppe Carleo, Garnet Kin-Lic Chan, Antonio Córcoles, Oliver Dial, Jay Gambetta, Artur Izmaylov, Caleb Johnson, Yukio Kawashima, David Kremer, Isaac Lauer, Peter Love, Guglielmo Mazzola, Takahito Nakajima, Paul Nation, Hanhee Paik, Emily Pritchett, Max Rossmannek, Paul Schweigert, James E.T. Smith, Miles Stoudenmire, Nobuyuki Yoshioka, Andrew Wack, and Christa Zoufal.
Data availability
All data supporting the findings are available from the corresponding authors upon reasonable request.
Supplementary Information: Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer
I Conventions and notation
I.1 Hamiltonian
Our starting point is the Born-Oppenheimer Hamiltonian, written in second quantization using a basis of $N_{\mathrm{MO}}$ orthonormal orbitals $\{\varphi_{p}\}_{p=1}^{N_{\mathrm{MO}}}$,
$\hat{H}=E_{\mathrm{0}}+\sum_{\begin{subarray}{c}pr\\ \sigma\end{subarray}}h_{pr}\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\phantom{\dagger}}_{r\sigma}+\sum_{\begin{subarray}{c}prqs\\ \sigma\tau\end{subarray}}\frac{(pr|qs)}{2}\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\dagger}_{q\tau}\hat{a}^{\phantom{\dagger}}_{s\tau}\hat{a}^{\phantom{\dagger}}_{r\sigma}\;.$ (S1)
@@ I.1 Hamiltonian
Indices $p,r,q,s$ label spatial orbitals in the basis set, $\sigma,\tau\in\{\uparrow,\downarrow\}$ are spin indices, and the fermionic operator $\hat{a}^{\dagger}_{p\sigma}/\hat{a}^{\phantom{\dagger}}_{p\sigma}$ creates/destroys an electron in orbital $p$ with spin $\sigma$. We define the number operator $\hat{n}_{p\sigma}=\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\phantom{\dagger}}_{p\sigma}$ which describes the number of electrons with spin $\sigma$ on orbital $p$. For non-relativistic all-electron calculations, the quantities
@@
I.2 Molecular species and active spaces
@@
The molecules simulated in this work are listed in Table 1. For N2 , we studied all non-core electrons and orbitals on quantum hardware. We computed the potential energy curve using (i) restricted Hartree-Fock (RHF) at 6-31G, and cc-pVDZ level [60, 61, 62] using PySCF [26, 63] and enforcing $D_{\infty h}$ symmetry and, after projecting the non-relativistic Born-Oppenheimer Hamiltonian in the space spanned by all non-core RHF orbitals with standard functionalities, with (ii) restricted and symmetry-preserving Moller-Plesset second-order perturbation theory (MP2), coupled cluster with singles and doubles (CCSD), complete active-space configuration interaction (CASCI), for 6-31G and selected configuration interaction, in its heat-bath flavor (HCI), for cc-pVDZ.
For [2Fe-2S] and [4Fe-4S] , we employed active spaces [4, 45], spanned by Fe[3d] and S[3p] orbitals, derived from a localized Density Functional Theory calculation with BP86 functional [64, 65], TZP-DKH basis [66], and sf-X2C (spin-free exact two-component) Hamiltonian [67, 68] to include scalar relativistic effects. We computed approximations to the ground-state energy with restricted RHF, MP2, CCSD, and HCI. The RHF, MP2, CCSD, CISD, CASCI and HCI calculations were carried out with the PySCF library [26].
I.3 Electron configurations and qubit mapping
@@
$|{\bf{x}}\rangle=\prod_{p\sigma}\left(\hat{a}^{\dagger}_{p\sigma}\right)^{x_{p\sigma}}|\emptyset\rangle,$ (S3)
@@
$|{\bf{x}}\rangle=\bigotimes_{p\sigma}|x_{p\sigma}\rangle,$ (S4)
@@
$|{\bf{x}}_{\mathrm{RHF}}\rangle=|\underbrace{0\dots 0}_{N_{\mathrm{MO}}-N_{\downarrow}}\underbrace{1\dots 1}_{N_{\downarrow}}\underbrace{0\dots 0}_{N_{\mathrm{MO}}-N_{\uparrow}}\underbrace{1\dots 1}_{N_{\uparrow}}\;\;\rangle,$ (S5)
@@
II Methods
II.1 Sample-based Quantum Diagonalization
This section provides a detailed description of the Sample-based quantum diagonalization (SQD) procedure. We describe the subspace projection and diagonalization, the self-consistent configuration recovery scheme, a lower bound to the likelihood of recovering different families of electronic configurations, and different techniques to optimize circuit parameters to minimize the estimator energy.
We first introduce some notations. We use $\mathcal{X}$ and $\tilde{\mathcal{X}}$ to denote a set of configurations sampled from probability distributions $P_{\Psi}({\bf{x}})=|\langle{\bf{x}}|\Psi\rangle|^{2}$ and $\tilde{P}_{\Psi}({\bf{x}})=\langle{\bf{x}}|\tilde{\rho}|{\bf{x}}\rangle$, where $\tilde{\rho}$ is a density operator corresponding to a noisy version of $|\Psi\rangle\langle\Psi|$. We denote $\mathcal{X}_{N}\subset\tilde{\mathcal{X}}$ the subset of configurations with the right particle number. The set of configurations recovered by the configuration recovery procedure is denoted by $\mathcal{X}_{\rightarrow N}$ and $\mathcal{X}_{\mathrm{R}}=\mathcal{X}_{N}\cup\mathcal{X}_{\rightarrow N}$ the set of configurations output by self-consistent configuration recovery, and $\mathcal{S}^{(k)}\subset\mathcal{X}_{\mathrm{R}}$ with $k=1\dots K$ a set of approximately $d$ configurations sampled from $\mathcal{X}_{\mathrm{R}}$. The wavefunction $|\psi^{(k)}\rangle$ in obtained by projection and diagonalization of $\hat{H}$ in the subspace spanned by the configurations in $\mathcal{S}^{(k)}$.
II.1.1 Eigenstate solver
Given a set of $d$ electronic configurations, the many-electron Hamiltonian is projected and diagonalized in the subspace spanned by the single-particle states defined by the electronic configurations, as proposed recently in [30]. We begin by considering a set of configurations $\mathcal{X}=\{{\bf{x}}_{(i)}\}$, all with the right-particle number for each spin sector. Sub-indices between parenthesis label configurations in a set and not configuration components. We use the generic label $\mathcal{X}$ for a set of configurations with the right particle number. In practice, the configuration recovery procedure will execute the eigenstate solver on the sets $\mathcal{X}_{N}$ and $\mathcal{X}_{\textrm{R}}$.

### v3

@@ Results
Figure 4: Experiments: Chemistry on large basis sets. (a) 58 qubits are used to model the N2 dissociation (cc-pVDZ basis set). (b) 45 qubits are used for the [2Fe-2S] cluster (TZP-DKH basis set) and (c) 77 qubits for the [4Fe-4S] cluster (TZP-DKH basis set). The top panels show a 3-dimensional representation of the geometry of each molecule. The middle panels show the qubits selected on a Heron quantum processor layout, following the same color convention as panel (b) in Fig. 3. The bottom panel in (a) shows the potential energy surface comparison, as well as the energy difference $\Delta E$ between the Heat-Bath Configuration Interaction (HCI) energy and the energies obtained from different methods. The brown scatter plot shows the value of $E^{(k)}$ for all batches of configurations and the connected dots show $\min_{k}(E^{(k)})$. The bottom panel in (b) shows the energy-variance analysis for three different eigenstates that both HCI and our method find upon increasing the value of $d$, as labeled by the color bar. For each approximate eigenstate $|\psi^{(k)}\rangle$, the horizontal axis $\Delta H=\langle\psi^{(k)}|\hat{H}^{2}|\psi^{(k)}\rangle-\langle\psi^{(k)}|\hat{H}|\psi^{(k)}\rangle^{2}$. The bottom panel in (c) shows a comparison of the energy-variance analysis applied to quantum measurement outcomes and bitstrings (with the correct particle number) sampled from the uniform distribution. The DMRG energy in panels (b) and (c) is from Ref.  (?).
[2Fe-2S] cluster: precision many-body physics
@@
Iron–sulfur (FeS) clusters are molecular ensembles of sulfide-linked 1- to 8-iron centers in variable oxidation states. They are important cofactors in biological processes ranging from nitrogen fixation to photosynthesis and respiration  (?). Their electronic structure, with multiple low-lying states of differing electronic and magnetic character, is responsible for their rich chemistry. At the same time, they pose considerable challenges for experimental studies and numerical tools. For our experiments we consider the synthetic [Fe2S2(SCH3)4]-2 cluster  (?), abbreviated [2Fe-2S] and used in numerical studies to mimic the oxidized dimers prominently found in ferredoxins  (?).
@@
The qubit mapping of the LUCJ circuit on the Heron processor for [2Fe-2S] is shown in the middle panel of Fig. 4 (b). We consider $K=10$ batches of configurations and $|\tilde{\mathcal{X}}|=2.4576\cdot 10^{6}$ measurement outcomes. The quantum runtime for this system is of $\sim 45$ minutes. For the [2Fe-2S] cluster, we perform an energy-variance analysis of the low-energy spectrum of the molecule. The energy-variance analysis is a tool routinely used in classical computational electronic structure to capture the convergence of the approximate eigenstate energy for different levels of accuracy of a computational method  (?). Here we use energy-variance analysis to assess the convergence as a function of $d$, which is directly related to quantum and classical accuracy, runtimes, and costs. If one can statistically sample from a good approximation of an eigenstate, points at finite number of samples will be distributed linearly in the energy-variance plane  (?). This gives us a tool to detect eigenstates for both quantum and classical methods.
[4Fe-4S] cluster: a stress test for methodology and quantum processors
@@
The circuits considered for the N2 experiments and [2Fe-2S] reached sizes of approximately 1-1.5k two-qubit gates. We now test the quality of the signal in noisy circuits that test the limits of Heron processors, using up to 6400 nodes of Fugaku for the classical processing. We consider the synthetic [Fe4S4(SCH3)4]-2 cluster  (?), abbreviated [4Fe-4S], a representative of nature’s cubanes, whose ground state deduction from experimental measurements was an early success of inorganic spectroscopy  (?). The LUCJ circuit employed for this molecular species contains approximately 3.5k two-qubit gates. The qubit mapping on the Heron processor for [4Fe-4S] is shown in the middle panel of Fig. 4 (c). As in the previous experiment, we consider $K=10$ batches of configurations and $|\tilde{\mathcal{X}}|=2.4576\cdot 10^{6}$ measurement outcomes. The quantum runtime for this system is of $\sim 45$ minutes. For $d>250\cdot 10^{3}$, configuration recovery is warm started with the ${\bf{n}}$ obtained from the method at $d=250\cdot 10^{3}$, and only two iterations are then performed.
@@
This last set of experiments sheds light on the quality of the quantum signal that is passed to the configuration recovery at these large circuit sizes. The bottom right panel in Fig. 4 (c) shows a comparison of the energy-variance analysis from measurement outcomes obtained from the Heron processor and configurations sampled from the uniform distribution. We see that, even if the quantum solutions produced are worse than other classical methods, the energy and variance obtained from quantum data are significantly lower than those obtained from uniformly distributed configurations (i.e. pure noise), on subspaces of the same size. This confirms that there is a valuable signal at circuit sizes of approximately 3.5k two-qubit gates.
Significance for quantum computing – Current quantum computers in isolation can perform calculations on systems sufficiently large that exact brute-force classical solutions are not available  (?, ?). However, these studies have targeted spin systems, leading to circuits that match the connectivity and the measurement and coherence budgets of the quantum devices. In this work, we present electronic structure calculations on active spaces beyond the scale where full configuration interactions are available. Key to achieve this result is the use of classical and quantum computers in concert, to implement the SQD method.
SQD makes economical use of quantum computing resources by drawing samples from a single quantum circuit. Although in principle other estimators, such as the standard ones used in variational quantum eigensolvers, have bounded variance for any wavefunction, the dire scaling of number of measurements to estimate energies makes them impractical for the molecules targeted in this work  (?). It is also more robust against quantum noise, because reconstructing the exact ground-state probability distribution on a quantum computer is not required to get accurate energy approximations, as long as one is sampling relevant configurations (i.e., in the ground-state support).
We have used a LUCJ class of quantum circuits that can reproduce a low-rank decomposition and sparsification of the quantum unitary CCSD (qUCCSD), which allowed us to keep circuit depths manageable  (?). Lower error rates on quantum operations will allow to access deeper quantum circuits with higher connectivity, giving access to more general probability distributions. We have performed optimization-free experiments exploiting the connection between LUCJ and classical coupled cluster theory, for the purpose of assessing accuracy and scalability, yet closing a quantum-classical optimization in future work will further improve the quality of our samples.
Generalization – The SQD method can be applied to simulation tasks other than quantum chemistry, if the target ground-state wavefunction can be accurately approximated by a sparse vector. Developing quantum circuits with polynomially-sized support in the computational basis will be an important element of the generalization of SQD, as these circuits are the sources of samples processed by classical computers.
@@
Implications in the search for quantum advantage – Our study contributes to the search for quantum advantage in at least two ways: by identifying an accuracy metric, and by formulating a set of conditions to improve over SCI.
Computations can be ranked against three parameters: runtime, energy or cost, and accuracy. While the first two are often easy to measure, ranking by accuracy is in general not straightforward. For methods that produce upper bounds to the ground-state energy, including SQD, the expectation value of the Hamiltonian defines an unconditional accuracy metric: a lower energy is ranked as a higher quality, all other conditions (e.g. total spin) being equal. Comparing SQD energies, for example, allowed us to benchmark our results against SCI and uniform configuration sampling on a 77-qubit experiment, without access to exact solutions. Additionally, the approximate wavefunctions produced here can be stored in classical memory, which permits a classical prover to certify them, and allows their manipulation by further classical processing.
@@
Using the expectation value of the Hamiltonian as an accuracy metric, one can easily and naturally rank SQD results along with those of variational classical methods, giving the search for this specific form of quantum advantage a quantitative meaning. Since every variational classical method has a specific domain of applicability  (?), identifying areas where SQD may offer an accuracy advantage is a delicate problem. For example, variational quantum Monte Carlo methods  (?, ?), of paramount importance in many-body physics, are sensitive to the structures of the probability distributions they are modeling, not just to their supports. Similarly, methods based on tensor networks  (?, ?) are successfully used to tackle strongly correlated problems in chemistry, granted the ability to converge their energies with bond dimensions, but convergence can be challenging in some cases  (?, ?) because of its computational cost and its sensitivity to the nature and ordering of the basis-set orbitals. Developing an in-depth understanding of SQD through extensive numeric and methodological investigations is necessary to establish or rule out advantage in strictly variational ground-state simulations.
Figure 5: Comparison of the quality of subspaces generated by HCI and an optimized LUCJ circuit in for the description of the ground state of [2Fe-2S]. Different panels show the energy as a function of subspace dimension of the corresponding diagonalization. The dots connected by arrows show the trajectory of HCI in the energy-subspace dimension plane. Each dot is labeled by the iteration it corresponds to. Different panels correspond to HCI calculations carried out with different values of the selection cutoff $\epsilon_{1}$, as indicated in the legend. Brown triangles show energies and subspace dimensions obtained after truncating the most accurate HCI wave function by repeatedly removing the electronic configurations corresponding to the lowest wave function amplitudes. The green triangles show the energies for different subspace dimensions obtained from the optimized LUCJ circuit. The green arrows show the subspace dimension that the circuit was optimized for (see main text). The horizontal solid and dashed lines indicate the DMRG estimates for the S0 and S1 eigenstates.
SQD shares with SCI the assumption that the ground state may be approximated by a sparse linear combination of determinants, i.e. with a number of determinants much smaller than the Hilbert space dimension. Note that this assumption does not necessarily underline variational Monte Carlo or tensor networks. Therefore, it is natural to look for conditions to improve over SCI. One such condition is the existence of a quantum circuit which produces subspaces of better quality, and more efficiently, than classical heuristic selection methods. In the search for ground states, the quality of a subspace can be determined by a lower variational energy. We conduct numerical experiments that suggest that there exists LUCJ circuits whose samples produce subspaces of better quality than the HCI classical selection heuristic. The LUCJ circuit under consideration shares the same depth and connectivity as the circuits used in the experiments. In this study we consider a particular flavor of SCI, Heat bath Configuration Interaction (HCI)  (?), and the [2Fe-2S] cluster (20 orbitals $\rightarrow$ 40+ qubits).
We first perform HCI calculations with different values of the selection cutoff $\epsilon_{1}$ (see Ref.  (?) for details on the definition of $\epsilon_{1}$), as shown in Fig. 5. We observe that the larger and more restrictive values of $\epsilon_{1}$ do not allow HCI to reach the DMRG energy reference. Instead, it converges to the first excited state (S1) energy. Decreasing the value of $\epsilon_{1}$ allows the subspace dimension to become larger and improve the energy beyond that first excited state, at the cost of a significant increase in subspace dimension, which results in a higher computational cost. From the largest HCI wave function (last point for $\epsilon_{1}=5\cdot 10^{-5}$), we repeatedly remove the electronic configurations whose wave function amplitude is below a given threshold and compute the energy in the resulting subspace, providing a collection of energy-subspace dimension pairs that we label as “upper bound to optimal” in Fig. 5. The removal of tens of millions of configurations (resulting in significantly smaller subspace dimensions) does not significantly deteriorate the quality of the ground-state approximation. These results indicates that HCI does not perform an optimal search of the relevant electronic configurations. Consequently, much smaller subspaces exist that yield comparable energy values. In this particular molecule, HCI needs to explore and perform diagonalizations in subspaces larger than the optimal search.
We optimize an LUCJ circuit to produce samples in the identified optimal subspaces with high probability. The optimal parameters are found in a two step optimization workflow. First, we optimize the Kullback–Leibler divergence between samples drawn from the LUCJ circuit and the amplitudes of 58 M-dimensional ground-state wave function estimation (green arrow in Fig. 5). Then, the circuit parameters are further fine-tuned to minimize the SQD energy using a differential-evolution strategy.
With the samples produced by the optimized LUCJ circuit, we proceed to evaluate the energy of resulting subspaces of varying dimensionality, including the 58 M-dimensional subspace that the circuit was optimized for. We observe good agreement between the LUCJ subspaces and those that are an upper bound to the optimal ones. It is worth noting that despite being optimized for a 58 M-dimensional subspace, the resulting circuit produces subspaces of outstanding quality of dimensionalities both smaller and larger.
The active space considered for the [2Fe-2S] cluster is small enough for HCI to be able to find the relevant bitstrings by exploring and performing diagonalizations in subspaces significantly larger than the optimal ones. This shows that in some problem instances, the diagonalization itself is not the runtime bottleneck. Instead, the runtime bottleneck is a sub-optimal proposal of electronic configurations. Larger and more strongly correlated systems will pose challenges for the classical heuristics. It is for these systems that one can look for a quantum advantage in sampling.
Characterizing the domain of applicability of different classical and quantum heuristics facilitates combining them in quantum-centric supercomputing environments, shifting the search for quantum advantage towards practical problems.
Materials and Methods
Conventions and notation
Hamiltonian
@@
Our starting point is the Born-Oppenheimer Hamiltonian, written in second quantization using a basis of $N_{\mathrm{MO}}$ orthonormal orbitals $\{\varphi_{p}\}_{p=1}^{N_{\mathrm{MO}}}$, as shown in Eq. 1. We define the number operator $\hat{n}_{p\sigma}=\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\phantom{\dagger}}_{p\sigma}$ which describes the number of electrons with spin $\sigma$ on orbital $p$. For non-relativistic all-electron calculations, the quantities
@@
Molecular species and active spaces
@@
The molecules simulated in this work are listed in Table 1. For N2 , we studied all non-core electrons and orbitals on quantum hardware. We computed the potential energy curve using (i) restricted Hartree-Fock (RHF) at 6-31G, and cc-pVDZ level  (?, ?, ?) using PySCF  (?, ?) and enforcing $D_{\infty h}$ symmetry and, after projecting the non-relativistic Born-Oppenheimer Hamiltonian in the space spanned by all non-core RHF orbitals with standard functionalities, with (ii) restricted and symmetry-preserving Moller-Plesset second-order perturbation theory (MP2), coupled cluster with singles and doubles (CCSD), complete active-space configuration interaction (CASCI), for 6-31G and selected configuration interaction, in its heat-bath flavor (HCI), for cc-pVDZ.
For [2Fe-2S] and [4Fe-4S] , we employed active spaces  (?, ?), spanned by Fe[3d] and S[3p] orbitals, derived from a localized Density Functional Theory calculation with BP86 functional  (?, ?), TZP-DKH basis  (?), and sf-X2C (spin-free exact two-component) Hamiltonian  (?, ?) to include scalar relativistic effects. We computed approximations to the ground-state energy with restricted RHF, MP2, CCSD, and HCI. The RHF, MP2, CCSD, CISD, CASCI and HCI calculations were carried out with the PySCF library  (?).
Electron configurations and qubit mapping
@@
$|{\bf{x}}\rangle=\prod_{p\sigma}\left(\hat{a}^{\dagger}_{p\sigma}\right)^{x_{p\sigma}}|\emptyset\rangle,$ (7)
@@
$|{\bf{x}}\rangle=\bigotimes_{p\sigma}|x_{p\sigma}\rangle,$ (8)
@@
$|{\bf{x}}_{\mathrm{RHF}}\rangle=|\underbrace{0\dots 0}_{N_{\mathrm{MO}}-N_{\downarrow}}\underbrace{1\dots 1}_{N_{\downarrow}}\underbrace{0\dots 0}_{N_{\mathrm{MO}}-N_{\uparrow}}\underbrace{1\dots 1}_{N_{\uparrow}}\;\;\rangle,$ (9)
@@
Sample-based Quantum Diagonalization
This section provides a detailed description of the Sample-based quantum diagonalization (SQD) procedure. We describe the subspace projection and diagonalization, and the self-consistent configuration recovery scheme. We also motivate the energy-variance analysis presented on Fig. 4.
In what follows, we use $\mathcal{X}$ and $\tilde{\mathcal{X}}$ to denote a set of configurations sampled from probability distributions $P_{\Psi}({\bf{x}})=|\langle{\bf{x}}|\Psi\rangle|^{2}$ and $\tilde{P}_{\Psi}({\bf{x}})=\langle{\bf{x}}|\tilde{\rho}|{\bf{x}}\rangle$, where $\tilde{\rho}$ is a density operator corresponding to a noisy counterpart of $|\Psi\rangle\langle\Psi|$. We denote $\mathcal{X}_{N}\subset\tilde{\mathcal{X}}$ the subset of configurations with the right particle number. The set of configurations recovered by the configuration recovery procedure is denoted by $\mathcal{X}_{\rightarrow N}$ and $\mathcal{X}_{\mathrm{R}}=\mathcal{X}_{N}\cup\mathcal{X}_{\rightarrow N}$ the set of configurations output by self-consistent configuration recovery, and $\mathcal{S}^{(k)}\subset\mathcal{X}_{\mathrm{R}}$ with $k=1\dots K$ a set of approximately $d$ configurations sampled from $\mathcal{X}_{\mathrm{R}}$. The wavefunction $|\psi^{(k)}\rangle$ in obtained by projection and diagonalization of $\hat{H}$ in the subspace spanned by the configurations in $\mathcal{S}^{(k)}$.
Eigenstate solver
Given a set of $d$ electronic configurations, the many-electron Hamiltonian is projected and diagonalized in the subspace spanned by the single-particle states defined by the electronic configurations, as proposed recently in  (?). We begin by considering a set of configurations $\mathcal{X}=\{{\bf{x}}_{(i)}\}$, all with the right-particle number for each spin sector. Sub-indices between parenthesis label configurations in a set and not configuration components. We use the generic label $\mathcal{X}$ for a set of configurations with the right particle number. In practice, the configuration recovery procedure will execute the eigenstate solver on the sets $\mathcal{X}_{N}$ and $\mathcal{X}_{\textrm{R}}$.